OpenStudy (anonymous):

64) The sum of the measures of the angles of a convex polygon is 900 degrees. Find the number of sides. a) 6 b) 7 c) 8 d) 9

OpenStudy (cwrw238):

sum interior angles = ( n - 2)* 180 so you need to solve 900 = (n - 2) * 180

OpenStudy (cwrw238):

can u continue?

OpenStudy (anonymous):

n=7

OpenStudy (cwrw238):

180n - 360 = 900 180n = 900 + 360 = 1260 n = 1260 / 180 right 7 is correct